OpenStudy (richyw):

The fourth corner of a parallelogram, S, is diagonally opposite P. Find S given P(1,1) Q(-3,4) R(6,10).

OpenStudy (richyw):

so why am I getting the prong answer when I add \[\vec{PQ}+\vec {PR}\]

OpenStudy (richyw):

wrong*

OpenStudy (anonymous):

show me how you add them

OpenStudy (richyw):

or I need to add this to the point P?

OpenStudy (richyw):

I add them to get (1,12). so I just need to add that vector to point P. how do I "show" that?

OpenStudy (richyw):

can I just say \[S=P+\vec{S}\]

OpenStudy (richyw):

does that make any since?

OpenStudy (richyw):

sense

OpenStudy (anonymous):

PQ=(-3-1,4-1)=(-4,3) PR=(6-1,10-1)=(5,9) PQ+PR=(1,12) now aply this vector to P to find S: (1,1)+(1,12)=(2,13)

OpenStudy (richyw):

ok thanks. that's exactly what I did. I was just unsure if you can just add a vector to a point. I mean I know you can (and you will get a point) but wasn't sure If I could just say that since math profs are picky about how you show things! thanks a lot

OpenStudy (anonymous):

yw